# Elasto-frictional reduced model of a cyclically sheared container filled with particles

### v1

@@ 5 Conclusion
In this paper, an elasto-hysteretic model describing a simply sheared granular medium confined in a box with rigid but moving walls has been studied using discrete element method. In particular, it was shown that a single degree-of-freedom Dahl model coupled with an elastic response reliably describes the behavior simulated by DEM, in terms of the stress-strain relation, the stored and the dissipated energies, ant the loss factor. The well-known effective medium theory predicting the elasticity of granular media, and a macroscopic effective friction coefficient, were identified as the main parameters of this model, via an extensive parametric study derived as a function of the confining pressure and the features of the particles and the mechanical excitation.
@@
Our study reveals that the elasto-hysteretic mechanical response of the granular medium mainly depends on a single dimensionless parameter, namely $\xi$, akin to a normalized Coulomb threshold. This parameter allows quantifying a priori how hysteretic versus how elastic is the mechanical response of our apparatus. The regime at high value of $\xi$ (i.e. high confinement pressure or low shear amplitudes) is asymptotically elastic, such that the loss factor vanishes in this regime. Nevertheless, the proposed model fails at low value of $\xi$ (i.e. low pressure or high shear amplitudes) due to large pressure fluctuation resulting from the anisotropy of the contact network. This effect leads to an additional nonlinearity of the elasticity, which ultimately results in the decrease of the damping ratio in this regime. Hence, in an intermediate range of $\xi$, the dissipation is dominated by an effective friction coefficient that is mostly independent of the particles features, whereas the stored energy is dominated by an effective elasticity that mostly depends on the confinement pressure and the elasticity of the particle. It thus results that the loss factor can be tuned by the confinement pressure, its maximum being independent of the particles and of the operating point.
@@
From the practical point of view, the reduced model derived in this study can be used in a more complex configurations, such as a granular medium clamped in between elastically bending beams to form a composite structure aiming at dissipating mechanical vibrations. Here, the advantage of a reduced model is to lower the complexity and the numerical cost of simulations coupling discrete element method with finite element method, hence allowing subsequent optimization process, for instance. The attempt to design a granular composite beam is a perspective under our consideration (Faulconnier, 2023). Also, the seminal features of this study has been patented (Fichant et al., 2024) by some of us in a separate work conducted in collaboration with Safran Spacecraft Propulsion, a major stakeholder of the Aeronautic Industry. The patented equipment aims at mitigating the vibrations of a cathode in the harsh environment of a spacecraft plasma engine, by using non-cohesive ceramic particles confined inside structural elements, as described in the present study.
@@
where $h$ corresponds to the contact overlap and $a$ to the particle diameter. In our study, the average contact strain ranges between $0.5\%$ and $4\%$, see Fig. A1.
@@
The effect of the nonlinearity resulting from the pressure fluctuation at low $\xi$ can be quantified by looking at the dependence of the Reynolds coefficient $R$ defined in Eq. 3 with the state variables $P$ or $\phi$ (both are related), and at the dependence of the nonlinear parameter $\rho=2\Delta P/3P$ defined in Eq. 20 with the pressure $P$ and the amplitude $\gamma_{m}$. Results are displayed on Fig. A2, showing well defined tendencies : (i) $R$ is an increasing function of a power law of the relative packing fraction $(\phi-\phi_{c})$; (ii) coherently, $R$ is an increasing function of a power law of the confining pressure $P$; (iii) $\rho$ is a decreasing function of $P$, meaning that loose packing leads to strong nonlinearity at a given amplitude $\gamma_{m}$; (iv) $\rho$ is an increasing function of $\gamma_{m}$, meaning that nonlinearity classicaly increases with the amplitude of perturbation.
@@
The fitting of the parameters $(\mu_{D},G_{D},G_{E})$ described in Secs. 3.3 and 3.5 is realized in two steps: the determination of an initial guess $(\mu_{D}^{\mathrm{ini}},G_{D}^{\mathrm{ini}},G_{E}^{\mathrm{ini}})$ by successive steps, followed by one refinement to obtain an optimal triplet $(\mu_{D}^{\mathrm{opt}},G_{D}^{\mathrm{opt}},G_{E}^{\mathrm{opt}})$. We first look for an initial ensemble $G_{D}^{\mathrm{ini}}(\mu_{D})$, as a function of a finite number of friction coefficient taken in a finite interval $0.01\leq\mu_{D}\leq 10$, which provides a hysteresis loop given by the model Eq. 4 having the same dissipated energy as in the DEM simulation. This step is made possible because, among the three parameters to fit, the dissipated energy does not depend on the conservative contribution associated with $G_{E}$. In practice, we determine $G_{D}^{\mathrm{ini}}$ at a given $\mu_{D}$ by minimizing (using Matlab’s simplex algorithm fminsearch) the root-mean-square error (RMSE) between the area within one stress-strain close loop estimated (using Matlab’s numerical integration trapz) by the model Eq. 4 and by the DEM simulation. This step provides a discrete numerical function, $G_{D}^{\mathrm{ini}}(\mu_{D})$, with which we then search for a value $G_{E}^{\mathrm{ini}}(\mu_{D})$ that minimizes the RMSE between the instantaneous stress response given by the model Eq. 4 and that of the simulation, at any given $\mu_{D}$. Finally, the initial guess $\mu_{D}^{\mathrm{ini}}$ is the value that provides the lowest aforementioned RMSE. The triplet ($\mu_{D}^{\mathrm{ini}},G_{D}^{\mathrm{ini}},G_{E}^{\mathrm{ini}}$) thus constitutes a reliable starting point to perform a 3-parameters minimization of the RMSE between the model Eq. 4 and the simulated instantaneous stress response, defined by the following cost function :
@@
This last step leads to a refined estimation of the optimal triplet $(\mu_{D}^{\mathrm{opt}},G_{D}^{\mathrm{opt}},G_{E}^{\mathrm{opt}})$. We provide in Fig. A3 the final value of the RMSE, defined in Eq. A2 and estimated at the optimal triplet, as a function of the dimensionless parameter $\xi$ for all the data presented in this study. The model Eq. 4 reliably fits in the range of $\xi\geq 0.1$ where the linear elasto-hysteretic regime prevails: the RMSE remains below $5$%. However, the fits provides poorer results at low $\xi$ where the nonlinear elastic contribution detailed in Sec. 4.5 becomes noticeable; in particular, the RMSE is larger than $10$% for the three data introduced in Sec. 4.5 (see the white triangles at lowest $\xi$), demonstrating the need to improve the model Eq. 4 as in Eq. 21.
@@
The authors declare that they have no conflicts of interest to this work. They do not have any commercial or associative interest that represents a conflict of interest in connection with the work submitted. The work described has not been submitted elsewhere for publication, in whole or in part, and all the authors listed have approved the manuscript that is enclosed.

### v2

@@ 5 Conclusion
In this paper, an elasto-hysteretic model describing a simply sheared granular medium confined in a box with rigid but moving walls has been studied using the discrete element method. In particular, it was shown that a single degree-of-freedom Dahl model coupled with an elastic response reliably describes the behavior simulated by DEM, in terms of the stress-strain relation, the stored and the dissipated energies, and the loss factor. The well-known effective medium theory predicting the elasticity of granular media, and a macroscopic effective friction coefficient, were identified as the main parameters of this model, via an extensive parametric study derived as a function of the confining pressure and the features of the particles and the mechanical excitation.
@@
Our study reveals that the elasto-hysteretic mechanical response of the granular medium mainly depends on a single dimensionless parameter, namely $\xi$, akin to a normalized Coulomb threshold. This parameter allows quantifying a priori the extent to which the mechanical response of our apparatus is hysteretic compared to elastic. The regime at a high value of $\xi$ (i.e. high confinement pressure or low shear amplitudes) is asymptotically elastic, such that the loss factor vanishes in this regime. Nevertheless, the proposed model fails at a low value of $\xi$ (i.e. low pressure or high shear amplitudes) due to large pressure fluctuation resulting from the anisotropy of the contact network. This effect leads to an additional nonlinearity of the elasticity, which ultimately results in the decrease of the damping ratio in this regime. Hence, in an intermediate range of $\xi$, the dissipation is dominated by an effective friction coefficient that is mostly independent of the particle features, whereas the stored energy is dominated by an effective elasticity that mostly depends on the confinement pressure and the elasticity of the particle. Thus the result is that the loss factor can be tuned by the confinement pressure, as its maximum is independent of the particles and operating point.
@@
From the practical point of view, the reduced model derived in this study can be used in more complex configurations, such as a granular medium clamped in between elastically bending beams to form a composite structure intended to dissipate mechanical vibrations. Here, the advantage of a reduced model is to lower the complexity and the numerical cost of simulations coupling the discrete element method with the finite element method, hence allowing subsequent optimization processes, for instance. We are now considering designing a granular composite beam (Faulconnier, 2023). Also, the seminal features of this study have been patented (Fichant et al., 2024) by some of our team in a separate work conducted in collaboration with Safran Spacecraft Propulsion, a major stakeholder of the Aeronautic Industry. The patented equipment is aimed at mitigating the vibrations of a cathode in the harsh environment of a spacecraft plasma engine, by using non-cohesive ceramic particles confined inside structural elements, as described in the present study.
@@
where $h$ corresponds to the contact overlap and $a$ to the particle diameter. In our study, the average contact strain ranges between $0.5\%$ and $4\%$, see Fig. A1. It is worth mentioning that the high confinement pressure considered for the nominal configuration, $P/k_{n}\simeq 0.2\%$, is chosen to accommodate the large shear deformation, $\gamma_{m}=0.15$, required to generate a clear hysteretic simulated response as in Fig. 1. It turns out that the typical normal stress within a contact is $(\sigma/k_{n})\propto(P/k_{n})^{1/3}\sim 0.1$ according to the Hertz potential Johnson (1985); very few materials can withstand such stress while remaining elastic, as for instance certain plastic materials and elastomers. However, the conclusions of the present study state that low shear amplitudes require low confinement pressure owing to the dependence of the mechanical response on the parameter $\xi\propto(P/k_{n})^{2/3}\gamma_{m}^{-1}$, see Eq. 16. For instance, glass particles were safely used in a patented prototype Fichant et al. (2024) designed to operate efficiently at moderate confinement pressures $P/k_{n}\ll 1$ in the range of vibration amplitudes $\gamma_{m}\ll 1$.
@@
The effect of the nonlinearity resulting from the pressure fluctuation at low $\xi$ can be quantified by looking at the dependence of the Reynolds coefficient $R$ defined in Eq. 3 with the state variables $P$ or $\phi$ (both are related), and at the dependence of the nonlinear parameter $\rho=2\Delta P/3P$ defined in Eq. 20 with the pressure $P$ and the amplitude $\gamma_{m}$. The results are displayed in Fig. A2, showing well defined trends: (i) $R$ is an increasing function of a power law of the relative packing fraction $(\phi-\phi_{c})$; (ii) also, $R$ is an increasing function of a power law of the confining pressure $P$; (iii) $\rho$ is a decreasing function of $P$, meaning that loose packing leads to strong nonlinearity at a given amplitude $\gamma_{m}$; (iv) $\rho$ is an increasing function of $\gamma_{m}$, meaning that nonlinearity increases classically with the amplitude of perturbation.
@@
The fitting of the parameters $(\mu_{D},G_{D},G_{E})$ described in Secs. 3.3 and 3.5 is performed in two steps: the determination of an initial guess $(\mu_{D}^{\mathrm{ini}},G_{D}^{\mathrm{ini}},G_{E}^{\mathrm{ini}})$ by successive steps, followed by one refinement to obtain an optimal triplet $(\mu_{D}^{\mathrm{opt}},G_{D}^{\mathrm{opt}},G_{E}^{\mathrm{opt}})$. We first look for an initial ensemble $G_{D}^{\mathrm{ini}}(\mu_{D})$ as a function of a finite number of friction coefficients taken in a finite interval $0.01\leq\mu_{D}\leq 10$, which provides a hysteresis loop given by the model Eq. 4 having the same dissipated energy as in the DEM simulation. This step is made possible because, among the three parameters to be fitted, the dissipated energy does not depend on the conservative contribution associated with $G_{E}$. In practice, we determine $G_{D}^{\mathrm{ini}}$ at a given $\mu_{D}$ by minimizing (using Matlab’s simplex algorithm fminsearch) the root-mean-square error (RMSE) between the area within one stress-strain close loop estimated (using Matlab’s numerical integration trapz) by the model Eq. 4 and by the DEM simulation. This step provides a discrete numerical function, $G_{D}^{\mathrm{ini}}(\mu_{D})$, with which we then search for a value $G_{E}^{\mathrm{ini}}(\mu_{D})$ that minimizes the RMSE between the instantaneous stress response given by the model Eq. 4 and that of the simulation, at any given $\mu_{D}$. Finally, the initial guess $\mu_{D}^{\mathrm{ini}}$ is the value that provides the lowest aforementioned RMSE. The triplet ($\mu_{D}^{\mathrm{ini}},G_{D}^{\mathrm{ini}},G_{E}^{\mathrm{ini}}$) thus constitutes a reliable starting point to perform a 3-parameter minimization of the RMSE between the model Eq. 4 and the simulated instantaneous stress response, defined by the following cost function:
@@
This last step leads to a refined estimation of the optimal triplet $(\mu_{D}^{\mathrm{opt}},G_{D}^{\mathrm{opt}},G_{E}^{\mathrm{opt}})$. We provide in Fig. A3 the final value of the RMSE, defined in Eq. A2 and estimated at the optimal triplet, as a function of the dimensionless parameter $\xi$ for all the data presented in this study. The model Eq. 4 fits reliably in the range of $\xi\geq 0.1$ where the linear elasto-hysteretic regime prevails: the RMSE remains below 55%. However, the fits provides poorer results at low $\xi$ where the nonlinear elastic contribution detailed in Sec. 4.5 becomes noticeable; in particular, the RMSE is larger than 1010% for the three data introduced in Sec. 4.5 (see the white triangles at the lowest $\xi$), demonstrating the need to improve the model Eq. 4 as in Eq. 21.
A4 Size and number of particles
Table A1: Parameters values for various combinations of particles’ number, size and polydispersity; the case (b) is the nominal configuration.
case $N_{p}$ $r_{p}$ ($m$) $dr_{p}/r_{p}$ ($\%$)
(a) 125 $(5\times 5\times 5)$ 2 25
(b) 1000 $(10\times 10\times 10)$ 1 25
(c) 8000 $(20\times 20\times 20)$ 0.5 25
(d) 1000 $(10\times 10\times 10)$ 1 40
This appendix provides details on the conditions at which the mechanical response of the representative elementary volume becomes independent on the number, the size and the polydispersity of particles, in order to get ride of the discrete and crystallized nature of the granular sample. The nominal configuration described in this study is a cubic volume containing $N_{p}=10^{3}$ particles with mean radius $r_{p}=1$ mm and $dr_{p}/r_{p}=25$% polydispersity. To justify this choice, we simulate and compare the stress-strain relation at the nominal packing fraction and shear amplitude $(\phi,\gamma_{m})=(0.70,0.15)$ for the few different configurations $(N_{p},r_{p},dr_{p}/r_{p})$ summarized in Tab. A1. Note that the representative elementary volume is kept constant in this parametric study, $N_{p}r_{p}^{3}=\mbox{const}$, to ensure that similar stress-strain curves rely on regimes with comparable energies. The four corresponding stress-strain responses are presented in Fig. A4. Looking first at the effect of the number of particles, the comparison of Figs. A4a and A4b shows that $N_{p}=5^{3}$ generates a non-smooth stress-strain relation as a hallmark of the discrete nature of sudden and local stick-slip rearrangements between a small number of frictional contacts, while eight times more particles, $N_{p}=10^{3}$, smooths the response by better averaging over more particles and events. However, the comparison between Figs. A4b and A4c shows that increasing the number of particles by another factor eight beyond the nominal configuration does not lead to further improvement. However, it is worth noting that this latter case is part of the dataset, see Tab. 1 and Figs. 456 and  7.
Figure A4: Shear stress-strain response for various cases of number of particles and polydispersity, over five cycles. The values of the parameters and the nomenclature used for the subfigures are listed in Tab. A1.
Looking then at the effect of the polydispersity, the quantitative and qualitative similarities of the responses shown in Figs. A4b and A4d indicates that increasing size randomness beyond the nominal case does not bring significant differences: we thus infer that the sample is already amorphous enough (i.e. its crystallization is reliably prevented) with a $25\%$ polydispersity. With a max-to-min size ratio of $1.25/0.75\simeq 1.66$, our size dispersion is indeed slightly larger than the bidispersity considered in Makse et al. (2004) (a distribution of grain radii in which $R_{1}=0.105$ mm for half the grains and $R_{2}=0.095$ mm for the other half; size ratio of $1.1$), the one in Xing et al. (2021) (an ensemble of beads with $5$ mm and $6$ mm diameters; size ratio of $1.2$), and the one in O’Hern et al. (2003) ($5050$ mixtures of particles with a size ratio of $1.4$), for the same purpose.
@@
The authors declare that they have no conflicts of interest with this work. They do not have any commercial or associative interest that represents a conflict of interest in connection with the work submitted. The work described has not been submitted elsewhere for publication, in whole or in part, and all the authors listed have approved the manuscript that is enclosed.